OpenStudy (henrietepurina):

Algebra 1 question Medal + Fan

OpenStudy (henrietepurina):

Create a system of equations that includes one linear equation and one quadratic equation. Part 1. Show all work to solving your system of equations algebraically. Part 2. Graph your system of equations, and show the solution graphically to verify your solution.

OpenStudy (henrietepurina):

can you hemp @cwrw238 ?

OpenStudy (henrietepurina):

*help*

OpenStudy (anonymous):

Do you have to make up your own system?

OpenStudy (henrietepurina):

Yeah

OpenStudy (cwrw238):

yes a linear equation could be y = x + 4 and a quadratic y = x^2 + 3x + 2

OpenStudy (henrietepurina):

Thank you, but is it a system?

OpenStudy (cwrw238):

yes the quadratic will give you a curved parabolic curve and the linear a straight line and will look something like